OpenStudy (marinebaby):

Can someone explain to me how to do this? I have no idea where to even start! Use dimensional analysis to find the number of cm/min in 5.8 m/hr.

OpenStudy (anonymous):

do you need cm/min or in/hr

OpenStudy (marinebaby):

I think cm/min!

OpenStudy (anonymous):



OpenStudy (anonymous):

What do you get for an answer?

OpenStudy (marinebaby):

Okay I understand it a little bit more but I think its supposed to be Meter converted to Centimeter! would it still be the same ?

OpenStudy (anonymous):

Oh shoot sorry...yeah you can just google the conversion for m to cm

OpenStudy (marinebaby):

it was 580 cm! What would I do now?

OpenStudy (marinebaby):

Thanks for the help by the way ill be sure to give you a medal! @Savannah_Lynn13 What would I do about the hour to minutes?

OpenStudy (hlilly2413):

Wait, what was your original question. Are you converting 560 mi/hr to cm/minute or the other way around?

OpenStudy (hlilly2413):

580*

OpenStudy (marinebaby):

its converting 5.8 meters / hour to centimeters / minute

OpenStudy (anonymous):

5.8 metres * 1 hour * 100 cm 1 hour * 60 minutes * 1 metres

OpenStudy (anonymous):

5.8 * 100 = 580 / 60= 9.67 cm/minute

OpenStudy (hlilly2413):